Argentina Partners is concerned about the possible effects of inflation on its operations. Presently, the company sells 75,000 units for $45 per unit. The variable production costs are $25, and fixed costs amount to $850,000. Production engineers have advised management that they expect unit labor costs to rise by 15 percent and unit materials costs to rise by 10 percent in the coming year. Of the $25 variable costs, 40 percent are from labor and 30 percent are from materials. Variable overhead costs are expected to increase by 20 percent. Sales prices cannot increase more than 10 percent. It is also expected that fixed costs will rise by 8 percent as a result of increased taxes and other miscellaneous fixed charges.

The company wishes to maintain the same level of profit in real dollar terms. It is expected that to accomplish this objective, profits must increase by 9 percent during the year.

Required:

a. Compute the volume in units and the dollar sales level necessary to maintain the present profit level, assuming that the maximum price increase is implemented. (Do not round intermediate calculations. Round up your answer for "Volume in units" to the nearest whole number and round your answer for "Sales" to the nearest whole dollar amount.)

b. Compute the volume of sales and the dollar sales level necessary to provide the 9 percent increase in profits, assuming that the maximum price increase is implemented. (Do not round intermediate calculations. Round up your answer for "Volume in units" to the nearest whole number and round your answer for "Sales" to the nearest whole dollar amount.)

c. If the volume of sales were to remain at 75,000 units, what price increase would be required to attain the 9 percent increase in profits? Calculate the new price. (Round your answer to 2 decimal places.)

My answers are wrong, i need the correct answers and how to find them ! my answers are

A.

Volume in units- 70282

Sales- 3478959

B-

Volume in Units -72905

Sales- 3608798

C.

48.88

ALL OF MY ANSWERS ARE WRONG SO PLEASE DO NOT PROVIDE THOSE ANSWERS !!
Answer #1

a. Computation of Volume in Units and Dollars sales required

Current Profit = Sales * (Sell price – Variable Cost) – Fixed Cost

Current Profit = 75000 * (45 – 25) – 850000

Current Profit = 1500000– 850000

Current Profit = $650000

New variable Cost = Change Material Cost + Changes Labor Cost + Changed Variable OH Cost

New variable Cost = 25 * 30% * 110% + 25 * 40% * 115% + 25 * 30% * 120%

New variable Cost = 8.25 + 11.50 + 9.00

New variable Cost = $28.75

Sale Price = Current Sale price * 110% = 45 * 110% = $49.5

New Fixed Costs = Current Fixed Costs * 108% = 850000 * 108% = $918000

Volume of Sales required = (Fixed Costs + Current profit) / (Sale price – Variable Cost)

Volume of Sales required = (918000 + 650000) / (49.5 – 28.75)

Volume of Sales required = 75566 Units

Dollar Sales required = Volume sales required * Sale Price

Dollar Sales required = 75566 * 49.5

Dollar Sales required = $3740517

b. volume of sales and the dollar sales level necessary to provide the 9 percent increase in profits, assuming that the maximum price increase is implemented

New Profit = Current Profit * 1.09 = 650000 * 1.09 = $708500

Volume of Sales required = (Fixed Costs + Current profit) / (Sale price – Variable Cost)

Volume of Sales required = (918000 + 708500) / (49.5 – 28.75)

Volume of Sales required = 78386 Units

Dollar Sales required = Volume sales required * Sale Price

Dollar Sales required = 78386 * 49.5

Dollar Sales required = $3880107

c. Price increase required to attain the 9 percent increase in profits & Computation of new price

Volume of Sales = (Fixed Costs + Current profit) / (Sale price – Variable Cost)

75000 = (918000 + 708500) / (Sale price – 28.75)

Sale price = 3782750 / 75000

New Sale price = $50.44

Increase sale price required = (New Price - Current price) / Current price

Increase sale price required = (50.44 - 45) / 45

Increase sale price required = 12.09%